OpenStudy (christos):

Probability, A test was developed to detect a particular type of arthritis in individuals over 50 years old. From a survey it is known that 10% of individuals in this age group suffer from this type of arthritis. The test was used on individuals confirmed with the disease, and a positive result was obtained 85% of the times. When the test was used on those confirmed without the disease, the negative result was reported 4% of the times. a) What is the probability that an individual has the disease given that the test is positive? b) What is the probability that an individual has the disease given that the test is negative?

OpenStudy (christos):

@ParthKohli @dan815 @Nnesha @nincompoop @zzr0ck3r @inkyvoyd @Somy @calculusxy

OpenStudy (christos):

@KyanTheDoodle @Skyz @aryandecoolest @Mchilds15 @Gokuporter @Ahsome @ayyookyndall

OpenStudy (adi3):

u know so many people

OpenStudy (adi3):

@Ahsome any idea

OpenStudy (christos):

:/

OpenStudy (adi3):

@Christos why did u make him ur fan if he is not replying

OpenStudy (adi3):

I thought he would be intelligent

OpenStudy (alekos):

I'd say it's 10% for both

OpenStudy (christos):

how did you solve that ? @alekos

OpenStudy (alekos):

The probability that an individual has the disease is 10% regardless of whether the test is positive or negative

OpenStudy (bookworm14):

but you have to think about what it is saying in the question specifically @phi

OpenStudy (phi):

we use P(a|b) = P(a and b)/P(b) to use this we need P( + test) P(+) = P(+|d) + P(+ / not d)

OpenStudy (christos):

hmm what is P( + test) ?

OpenStudy (phi):

Probability of getting a positive test if you have the disease it is 0.85 if you don't have it it is 0.04 (i.e. the test responds postively, even though there is no disease)

OpenStudy (christos):

so the answer to a is 85% ?

OpenStudy (christos):

in this situation d is 0.1 ?

OpenStudy (phi):

my thinking might be muddled here, and it will take a few minutes to clear it up

OpenStudy (christos):

hmm ok

OpenStudy (phi):

Ok, here are the gory details Some theorems 1 Pr(a | b) = Pr (a & b) / Pr(b) from which we get 2 Pr(a & b) = Pr(b) * Pr(a | b) 3 Pr(a) = Pr(b) * Pr(a|b) + Pr(~b) * Pr(a | ~b) Given Pr(d)= 0.1 Pr(~d) = 0.9 Pr(+ | d) = 0.85 Pr(+ | ~d) = 0.04

OpenStudy (phi):

we want to find Pr(d|+) which by eq 1 is Pr(d | +) = Pr( d & +) / Pr(+) which means we have to find both Pr's on the right side because they are not given to us explicitly

OpenStudy (phi):

By eq 2. Pr(a & b) = Pr(b) * Pr(a | b) Pr (d & +)= Pr( + & d) and Pr( + & d) = Pr(d) * Pr ( + | d) now we have info for this: Pr(d)= 0.1, Pr(+|d) =0.85 so Pr(+ & d) = 0.085

OpenStudy (phi):

btw I am using Pr( + & d) to mean the prob of the intersection of positive result and having the disease.

OpenStudy (phi):

so far we have Pr(d | +) = Pr( d & +) / Pr(+) or Pr(d | +) = 0.085 / Pr(+) now we need Pr(+) we use eq 3 Pr(a) = Pr(b) * Pr(a|b) + Pr(~b) * Pr(a | ~b) Pr(+) = Pr(d) * Pr(+|d) + Pr(~d) * Pr(+ | ~d) we have these numbers, so Pr(+) = 0.1 * 0.85 + 0.9 * 0.04 Pr(+)= 0.121 we get Pr(d | +) = 0.085 / 0.121 = 0.702 or about 70.2%